How do you write an equation for a parabola opening upward with x intercepts -3 and 4?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How do you write an equation for a parabola opening upward with x intercepts -3 and 4?
-----------
y = (x+3)*(x-4) is one way.
===========
y = k*(x+3)*(x-4) where k is any real number > 0.


Answer by stanbon(75887)

How do you write an equation for a parabola opening upward with x intercepts -3 and 4?
-------------------
y = (x+3)(x-4)
--------------------
Comment: Opens upward because x^2 has a positive coefficient.
----------------------
Cheers,
Stan H.